LONG TERM STABILITY IN CW-CRDS EXPERIMENTS

\underline{HAIFENG HUANG}, ; KEVIN K. LEHMANN, Department of Chemistry, University of Virginia, Charlottesville VA, 22904-4319.

Allan variance can be used to characterize the slow drift of CW cavity ring-down spectroscopy (CW-CRDS) system, caused by the instability of experimental environments. By including the differential measurement of on- and off-peak decay rates in the system, the drifting of both decay rates can cancel out greatly. A sensitivity of 5.6 × 10^{−12} cm^{−1} during an optimum integration time of 15.4 minutes has been obtained with our CW-CRDS system. This sensitivity corresponds a 3σ methane detection limit of 0.3 parts per billion by volume (ppbv) in N_{2} at 20 torr or 37 parts per trillion by volume (pptv) in N_{2} at 760 torr measurement pressure, near 1.65 μm laser wavelength.